OpenStudy (anonymous):

Percent of Change Original 50 New 75 Help!!!!!

OpenStudy (anonymous):

50*.5=25 50*1.5=75 the answer is 1.5

OpenStudy (anonymous):

the formula for percent of change is (y2-y1/y1)*100 75-50/50= 25/50=1/2=.5*100=50%

OpenStudy (anonymous):

THANK YOU AWESOME PEOPLE :D

OpenStudy (anonymous):

:D